OpenStudy (anonymous):

a line is prepindicular to the line y=2x+3 and has the same x-intercept as x+3y+10=0. find the equation for this line in the y=mx+b form!

OpenStudy (anonymous):

the slope is -1/2 the x int is (-10, 0)

OpenStudy (anonymous):

so y/ (x+10) = -1/2 2y= -10-x y=-x/2 -5

OpenStudy (anonymous):

since it is perpendicular to y=2x+3 , slope =-1/2, then x-intercept for the second line is -10. therefore the line is y=-1/2*(x+10) y=-x/2 -5

OpenStudy (amistre64):

ugh.. y = (-1/2)(x+10)+0 right?

OpenStudy (anonymous):

thnx to you all:D